A solid cylinder of radius r(1)=2.5cm length L(1)=5.0cm and temperature 40°C is suspended in an environment of temperature 60°C . The thermal radiation transfer rate for cylinder is 1.0Watt. if the cylinder is stretched until its radius becomes r(2)=0.50cm the thermal radiation transfer rate is changed to :
Ans 3.35w
1
Expert's answer
2020-10-30T13:18:28-0400

r1 = 0.025 m

T = 40 + 273 = 313 K

L1 = 0.05 m

Tenv = 60 + 273 = 333 K

P1 = 1.0 Watt

r2 = 0.005 m

P1 = σεA1(Tenv4 – T4)

P2 = σεA2(Tenv4 – T4)

"σ = 5.6704 \times 10^{-8} \; \frac{W}{m^2K^4}" Stefan Boltzmann constant

"A_1 = \u03c0r_1^2 + \u03c0r_1^2 +(2 \u03c0r_1)L_1"

"A_1 = 3.14 \times 0.025^2 + 3.14 \times 0.025^2 +(2 \times 3.14 \times 0.025) \times 0.05 = 0.011780972 \;m^2"

"ε = \frac{P_1}{σA_1(T_{env}^4 – T^4)}"

"ε = \frac{1}{(5.6704 \times 10^{-8})(0.011780972)(333^4 – 313^4)} = 0.555"

The stretched cylinder and the former have the same volume:

V1 =V2

πr12L1 = πr22L2

"L_2 = \frac{πr_1^2L_1}{πr_2^2}"

"L_2 = \frac{0.025^2 \times 0.05}{0.005^2} = 1.25 \;m"

"A_2 = \u03c0r_2^2 + \u03c0r_2^2 +(2 \u03c0r_2)L_2"

"A_2 = 3.14 \times (0.005)^2 + 3.14 \times (0.005)^2 + (2\times 3.14 \times 0.005) \times 1.25 = 0.039407 \;m^2"

P2 = σεA2(Tenv4 – T4)

"P_2 = (5.6704 \times 10^{-8})(0.555)(0.039407)(333^4 – 313^4) = 3.346\; W"